### v2

# TXL Fusion: A Hybrid Machine Learning Framework Integrating Chemical Heuristics and Large Language Models for Topological Materials Discovery

Abstract
Topological materials—including insulators (TIs) and semimetals (TSMs)—hold immense promise for quantum technologies, yet their discovery remains constrained by the high computational cost of first-principles calculations and the slow, resource-intensive nature of experimental synthesis. Here, we introduce TXL Fusion, a hybrid machine learning framework that integrates chemical heuristics, engineered physical descriptors, and large language model (LLM) embeddings to accelerate the discovery of topological materials. By incorporating features such as space group symmetry, valence electron configurations, and composition-derived metrics, TXL Fusion classifies materials across trivial, TSM, and TI categories with improved accuracy and generalization compared to conventional approaches. The framework successfully identified new candidates, with representative cases further validated through density functional theory (DFT), confirming its predictive robustness. By uniting data-driven learning with chemical intuition, TXL Fusion enables rapid and interpretable exploration of complex materials spaces, establishing a scalable paradigm for the intelligent discovery of next-generation topological and quantum materials.
keywords:
American Chemical Society,
[…]
Topological materials, encompassing topological insulators (TIs)hasan2010topological, qi2011topological and topological semimetals (TSMs),liu2014discovery, xu2015discovery represent unconventional quantum phases of matter characterized by nontrivial electronic band topology. Their robust boundary states, protected against perturbations such as disorder or symmetry breaking, give rise to exotic phenomena including the quantum spin Hall effect,kane2005quantum unusual transport properties,qi2011topological and magnetoelectric responseshu2019transport. These unique properties position topological materials as promising candidates for next-generation quantum and spintronic technologies. Since the emergence of the field, a central challenge has been the reliable identification and classification of such materials. Early efforts relied heavily on first-principles calculations combined with topological band theory,xiao2021first, bansil2016band a computationally intensive but powerful route for establishing topological order. The advent of symmetry indicatorsslager2013space, kruthoff2017topological, po2017symmetry and topological quantum chemistrybradlyn2017topological represented a major milestone, enabling efficient diagnosis of many topological phases directly from symmetry representations of electronic states. These symmetry-based approaches facilitated high-throughput computational searches, resulting in extensive databases of candidate topological materials and accelerating both theoretical and experimental exploration.vergniory2019complete, zhang2019catalogue, tang2019comprehensive
Despite these advances, symmetry-based methods face inherent limitations. Certain topological phases, such as Chern insulators and time-reversal-invariant $Z_{2}$ insulators without point group symmetries, remain invisible to symmetry indicators and require explicit evaluation of wavefunction-based topological invariants, which is computationally expensive.po2017symmetry Materials with low-symmetry or complex magnetic structures pose additional challenges for symmetry-based diagnosis.xu2020high As a result, the discovery of topological materials is constrained by computational bottlenecks and the limited scope of existing frameworks.
Over the past decade, machine learning (ML) has become a scalable alternative to symmetry-based approaches for classification and properties prediction of topological materials.zhang2018machine, scheurer2020unsupervised, cao2020artificial, choudhary2021high, schleder2021machine, tyner2024machine, hong2025discovery, haosheng2025predicting Models such as gradient-boosted trees trained on space group (SG), electron count, and orbital-resolved valence descriptors have achieved strong performance,claussen2020detection and neural networks applied to computed XANES spectra have further expanded predictive capabilities.nina2022machine Despite these advances, conventional ML models operate solely on structured numerical inputs, limiting their ability to incorporate unstructured information—such as material descriptions, experimental annotations, or insights from scientific literature. To overcome data and scalability constraints, composition-based heuristics have been proposed, most notably the topogivity score $g(M)$,ma2023topogivity with subsequent extensions such as the integration of Hubbard $U$ parameter for magnetic systemsxu2024discovering and the inclusion of quantum correlations across elementsxu2025quantum. While efficient and interpretable, these composition-only rules remain insensitive to essential physical features and often struggle to distinguish closely related phases, particularly TSMs and TIs.
Recently, large language models (LLMs) have opened new opportunities by encoding chemical knowledge from vast scientific corpora. Unlike conventional descriptor-based ML, LLMs capture contextual relationships and support few-shot learning without manual feature engineering. Their versatility spans fine-tuning for chemistry Q & A,Jablonka2024 hybrid embedding with graph neural networks,li2025hybrid synthesis prediction,kim2024llm chatbot-assisted quantum chemistry,gadde2025chatbot dataset curation,kang2025harnessing AI-driven simulation assistants,hu2025aitomia and crystalline property prediction with transformer architectures.rubungo2025llm, korolev2023accurate Beyond prediction, LLM embeddings provide compositional and structural representations that enable similarity search, candidate retrieval, and multi-task learning,qu2024llm while fine-tuning on text-encoded atomistic data has shown promise in generating physically stable structures.gruver2024llm
Despite recent advances, the use of LLMs in the discovery and categorization of topological materials remains largely underexplored. To bridge this gap, we introduce TXL Fusion, a hybrid framework that unites three complementary pillars: (i) composition-driven chemical heuristics, (ii) domain-specific numerical descriptors, and (iii) embeddings derived from a fine-tuned LLM. By leveraging the strengths of these distinct sources, TXL Fusion delivers higher accuracy, robustness, and generalization than any single method alone. The framework enables high-throughput screening of unexplored chemical spaces, identifying numerous potential topological materials, with a subset of low-cost cases further validated through density functional theory (DFT) calculations. Our results demonstrate that combining symbolic, statistical, and linguistic knowledge provides a powerful paradigm for addressing complex discovery challenges in materials science.
1 Dataset and Feature Selection
We source our data from the topological materials database,topo_materials, bilbao_cryst, bradlyn2017topological, vergniory2019complete, vergniory2022all which includes DFT calculations with spin–orbit coupling (SOC) on 38,184 materials, comprising 6,109 TIs ($\sim$16%), 13,985 TSMs ($\sim$36.6%), and 18,090 trivial materials ($\sim$47.3%). Guided by both theoretical considerations and systematic empirical analysis, we conduct a comprehensive feature selection process where our initial feature set spanned over many properties including chemical bonding characteristics (e.g., covalent vs. ionic tendencies), SOC strength ($\propto Z^{4}$), periodic table’s group and column positions, total number of electrons, SG, valence electrons and atomic mass. Through iterative evaluation, we refine this broad feature pool to a compact set of descriptors that consistently offered both statistical robustness and physical interpretability. Further methodological details and extended analysis are provided in Section S1 of Supporting Information. Based on our analyses, SG symmetry emerged as the most decisive indicator of topological character. High-symmetry cubic and tetragonal SGs (e.g., 194, 225, 221, 139) are predominantly associated with TSMs, while low-symmetry monoclinic and orthorhombic SGs (e.g., 14, 62, 15) favor trivial compounds. TIs occupy intermediate symmetry regimes (e.g., 62, 63, 139), indicating that symmetry constraints are necessary but not sufficient for topological behavior. Several SGs are entirely absent in specific classes, confirming strong symmetry selectivity across topological phases (Supporting Information, Fig. S1 and Table S1).
Complementary chemical and electronic descriptors further enhance class separability. TIs show enriched d– and f–orbital participation and higher transition-metal and lanthanide content, consistent with strong SOC and band inversion. Trivial compounds, by contrast, are dominated by nonmetals and p–orbital bonding, reflecting localized, covalent environments. Electron-count parity further differentiates metallic from insulating systems: 70.7% of TSMs possess odd electron counts, enforcing metallicity via Kramers degeneracy, while most TIs and trivials exhibit even counts that permit full band filling. Bonding analysis reveals that TIs and TSMs preferentially adopt mostly covalent character, whereas trivials are more ionic, underscoring the role of delocalized orbitals in stabilizing nontrivial topology (Supporting Information, Table S2).
Collectively, these insights establish a concise and interpretable feature space—integrating symmetry, orbital, compositional, and bonding descriptors—that forms the foundation of our TXL Fusion framework presented below.
Figure 1: Schematic flowchart of the TXL Fusion model, outlining the main stages of the workflow.
2 TXL Fusion architecture
The TXL Fusion model integrates chemically inspired heuristics, numerical descriptors, and LLM embeddings within a unified hybrid framework that couples domain intuition with data-driven learning for robust classification of topological materials. As illustrated in Fig. 1, the framework consists of three interconnected modules, each detailed in the Supporting Information (Section S2).
The pipeline begins with a composition-based heuristic module, adapted from Ma et al.ma2023topogivity, which assigns elemental contribution scores to estimate the likelihood of a material being trivial, TSM, or a TI. This heuristic captures global compositional trends consistent with chemical intuition—where lighter, nonmetallic elements favor trivial phases, while heavier elements such as Bi, Sb, and Te correlate with topological behavior. However, since heuristic trends alone cannot distinguish between TIs and TSMs (see Table 1 in the "Results and Discussion" section), we complement it by numerical descriptor module.
The numerical descriptor module encodes physically meaningful quantities—such as SG symmetry, total and parity-resolved electron counts, orbital occupancies, electronegativity differences, and compositional ratios—into a fixed-length vector. These descriptors, selected through the feature analysis described in Section "Dataset and Feature Selection" and Supporting Information (Section S1), provide a systematic and interpretable representation of the material’s underlying physics.
The third component in our pipeline is LLM embedding module, built upon a fine-tuned SciBERT encoder, converts structured textual descriptions of materials (including chemical formulas, SG annotations, orbital contributions, and heuristic-derived reasoning) into dense semantic embeddings. These embeddings capture contextual and higher-order correlations beyond what explicit numerical features represent, linking symbolic chemical knowledge with statistical learning.
Finally, the heuristic outputs, numerical descriptors, and LLM embeddings are concatenated to form a comprehensive feature representation, which is passed to an eXtreme Gradient Boosting (XGB) classifier for final prediction. This multi-layered integration enables TXL Fusion to balance interpretability and performance, offering a scalable, generalizable approach for the intelligent discovery of topological materials. Detailed implementation procedures and model specifications are provided in the Supporting Information (Sections S2-3).
3 Results and Discussion
In this section, we assess the capability of the proposed TXL Fusion model to distinguish topological materials from trivial ones and benchmark it against two standalone baselines: the composition-based heuristic rule $g(M)$ and a numerical descriptor-based XGB model. Detailed training procedures are provided in the Supporting Information (Section S3). The dataset was divided into 80% for training and 20% for testing (7637 materials), with the latter designated as Discovery Space-1, on which high-throughput screening was conducted using features learned from the training data. The training subset was further partitioned into 80% sub-training and 20% validation data; no validation results are reported for $g(M)$ since it represents a fixed analytical rule. Model performance is evaluated using precision, recall, and F1-score:
$\text{Precision}=\frac{TP}{TP+FP},\quad\text{Recall}=\frac{TP}{TP+FN},\quad\text{F1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}},$ (1)
where $TP$, FPFP, and $FN$ denote true positives, false positives, and false negatives. Precision reflects the fraction of predicted positives that are correct, recall measures the fraction of true positives recovered, and F1-score balances the two.
Table 1 summarizes comparative performance on 7637 unseen materials of Discovery Space-1. The $g(M)$ rule, while interpretable and chemically motivated, performs poorly in Discovery Space-1: although $g(M)^{\mathrm{Trivial-Others}}$ distinguishes trivial compounds reasonably well (F1 = 0.81), it performs worse on TSMs (F1 = 0.62) and fails entirely for TIs (F1 = 0.00), reflecting its reliance on composition vectors alone. The numerical descriptor based XGB model, leveraging richer numerical and structural features, outperforms $g(M)$ rule for TI and TSM classification (F1 > 0.85 across splits) but still struggles with TIs (F1 = 0.57 in Discovery Space-1), limited by its factorized nature, which treats features such as SG, electron count, orbital occupations, and elemental ratios as independent scalars.
By contrast, TXL Fusion consistently outperforms the baselines across all classes, achieving F1 scores of 0.89 (trivial), 0.89 (TSM), and 0.62 (TI). This represents a clear improvement over XGB, particularly for TIs by +5%. This superior performance stems from its hybrid design, with its primary advantage arising from LLM embeddings that capture complex, coupled electronic–structural correlations, allowing the model to surpass the limitations of factorized numerical descriptors. A detailed comparison of semantic versus numerical descriptors, highlighting these advantages, is provided in Section S4 of the Supporting Information.
It is worth mentioning that performance varies with chemical complexity (Supporting Information Section S5). For single-element compounds, both TXL and XGB struggle with TIs (F1 < 0.25) due to data scarcity, limited descriptor variability, and subtle, non-compositional physical effects. Accuracy improves for 2–3 element compounds, where TXL slightly outperforms XGB (F1 $\approx$ 0.62–0.64 vs. 0.59–0.63). The largest advantage appears for 4–5 element systems, where TXL sustains reliable TI predictions (F1 = 0.57–0.61) while XGB sharply underperforms (F1 = 0.24–0.31). For six-element compounds—the most sparsely represented category in the dataset—TXL Fusion retains better balance for rare TI cases (F1 = 0.33 compared to 0.00 for XGB), while both models achieve strong performance for trivial and TSM classes.
Calibration results (Supporting Information Section S5) show that TXL Fusion is most reliable at higher confidence levels: predictions above 90% confidence are nearly always correct across all classes, supported by sufficient sample counts. In contrast, predictions in the 30–50% range are less dependable, with accuracies often below bin midpoints, indicating overconfidence. This is most evident in the 30–40% range, where trivial predictions show perfect accuracy but are based on only three samples, rendering the result statistically insignificant.
Finally, feature importance analysis (Fig. 2) highlights TXL Fusion’s advantage. While the numerical descriptor-based XGB model depends heavily on electron parity, SG probabilities, and p-valence counts, TXL Fusion derives balanced contributions from $g(M)$ scores and LLM embeddings. This integration of interpretability, engineered robustness, and contextual depth underlies TXL Fusion’s superior generalization in discovery-oriented test spaces.
Figure 2: Feature importance in material classification for (A) the numerical descriptor-based XGB model and (B) the TXL Fusion model. Here Bert_n ($n=1,\dots,5$) denotes the five principal components derived from principal component analysis (PCA) of the 768-dimensional embeddings obtained from the fine-tuned LLM.
[…]
Model Class Validation Set Discovery Space-1
Precision Recall F1 Precision Recall F1
$g(M)^{\rm{Trivial-Others}}$ Trivial 0.87 0.77 0.81
Others 0.77 0.87 0.81
$g(M)^{\rm{TSM-Others}}$ TSM 0.65 0.60 0.62
Others 0.78 0.81 0.80
$g(M)^{\rm{TI-Others}}$ TI 0.00 0.00 0.00
Others 0.83 1.00 0.90
XGB model Trivial 0.88 0.91 0.89 0.87 0.90 0.88
TSM 0.84 0.89 0.86 0.82 0.88 0.85
TI 0.66 0.49 0.56 0.67 0.50 0.57
TXL Fusion Trivial 0.90 0.91 0.91 0.88 0.91 0.89
TSM 0.88 0.87 0.89 0.85 0.87 0.86
TI 0.68 0.63 0.65 0.67 0.58 0.62
To further evaluate the TXL Fusion’s predictive capability, we used the set of 1,433 materials reported in Ma et al.ma2023topogivity as a discovery space. These materials, originally curated by Tang et al.tang2019comprehensive, are of particular interest because their topological character cannot be resolved using symmetry indicators. Among them, 1,235 compounds are already present in the topological materials database, leaving 198 as a new testing set, which we refer to as Discovery Space-2. Two compounds belonging to SG 178 were excluded, as this SG is not represented in the topological materials database, resulting in 196 candidate materials.
From these 196 candidates, our TXL Fusion model identified 21 potential TSMs with varying confidence levels: Li22Pb5 (216), Bi2Cl7Se5 (19), Cl11Mo3N2 (29), Li22Sn5 (216), AgPb4Pd6 (152), NS2 (102), SbO2 (33), P3Rb2Se6 (29), RbSO3 (150), LaMo2O5 (186), TlC2O2 (19), Ta21Te13 (183), OTi6 (159), CNSe (19), RbGe8Li7 (186), Ag10Br3Te4 (36), Cs9O3Tl4 (197), C8Cs (191), In11Mo40O62 (26), BiSe3Sr (19), Ge5Li22 (216), P3Sc7 (186).
Among these, In11Mo40O62 (26) and AgPb4Pd6 (152) were flagged as TSMs but could not be located in the materials project databasehorton2025matproj and were thus excluded. For the remaining 19 predicted TSMs, DFT calculations were carried out on five representative compounds (CsC8 (191), OTi6 (159), SbO2 (33), NS2 (102), and P3Sc7 (186)), while the others were omitted due to the high computational cost (see Supporting Information Section S6 for computational details).
As shown in Fig. 3, among the five tested compounds—four (CsC8, OTi6, SbO2, and P3Sc7) were validated as TSMs, suggesting an estimated success rate of approximately 80% if extrapolated to the full set. Specifically, as shown in Fig. 3, CsC8 (191) exhibits graphite-derived dispersions with weak SOC, maintaining a trivial character, while OTi6 (159) displays broad metallic states without inversion. SbO2 (33), despite containing the heavy pnictogen Sb, shows complex near-metallic behavior with small gaps and pseudogaps across the Brillouin zone. P3Sc7 (186) features dispersive metallic bands consistent with trivial metallicity. These results demonstrate that the TXL Fusion model accurately captures the topological nature of materials directly from elemental and compositional features, highlighting its strong predictive capability and potential as a scalable framework for data-driven topological materials discovery.
Figure 3: Electronic band structures along with space groups (A) CsC8 (191), (B) OTi6 (159), (C) SbO2 (33) and (D) P3Sc7 (186).
4 Concluding remarks
In this work, we introduced TXL Fusion, a hybrid framework that integrates chemically informed heuristics, numerical descriptors, and LLM-derived semantic embeddings for robust classification of topological materials. Across all classes, TXL Fusion consistently surpasses both the standalone XGB model and the heuristic $g(M)$ rule, demonstrating the effectiveness of combining domain intuition with data-driven, context-aware representations. By unifying interpretable chemical trends with high-dimensional semantic embeddings, the model achieves improved accuracy and generalization, particularly for compositionally complex systems. To promote broader community access, TXL Fusion will be made publicly available through the Aitomistic Hub at aitomistic.xyz.
Despite these advances, several intrinsic challenges remain. TIs are the most difficult class to predict, primarily due to label uncertainty and data sparsity. As discussed in Section S7 of the Supporting Information, DFT-based topological labels—especially for small-gap or borderline systems—can be unreliable, effectively introducing noise into the training data. Consistent with this interpretation, evaluation on high-confidence TI subsets shows systematic performance improvements as label reliability increases (Fig. S7). In addition, element-resolved analysis in Section S5 reveals reduced performance for low-complexity compounds (single- and binary-element systems), arising from a combination of data scarcity, limited descriptor expressiveness due to compositional features collapsing to constants or narrow ranges, and the dominance of subtle, non-compositional physical effects (see Section S5.1).
While these limitations could be mitigated by incorporating more localized or symmetry-resolved physical descriptors—such as symmetry-indicator eigenvalues, band inversion indicators, Wannier charge centers, or orbital-projected band-edge features—such quantities are computationally intensive and not uniformly available at scale. Within our framework, we therefore prioritize broad chemical coverage, enabling large-scale screening based solely on elemental information without requiring prior expensive calculations.
At the same time, TXL Fusion demonstrates strong generalization under chemical sparsity. As detailed in Section S5.2, the model maintains reliable performance for trivial and TSM phases even in sparsely sampled or partially out-of-distribution chemical families, with its advantage over numerical baselines increasing alongside chemical and electronic complexity. Moreover, confidence-based calibration analysis (Section S5) shows that TXL Fusion is well calibrated at high confidence levels, indicating that its most confident predictions are highly reliable and well suited for guiding experimental validation.
More broadly, this study highlights the power of LLM-based semantic embeddings for materials discovery. As discussed in Section S4, these embeddings capture coupled electronic–structural relationships that are fragmented in conventional factorized numerical descriptors, enabling improved discrimination of subtle topological regimes. Looking forward, continued improvements in label fidelity, expanded high-confidence datasets, and the incorporation of additional physics-aware learning strategies—such as transfer learning across chemical families or cost-sensitive training to address class imbalance—are expected to further enhance performance. More generally, the TXL Fusion framework is not limited to topological materials and provides a flexible foundation for accelerating discovery across a broad range of quantum and functional materials systems.
{acknowledgement}
A.U. acknowledges funding from the National Natural Science Foundation of China (No. W2433037) and the Natural Science Foundation of Anhui Province (No. 2408085QA002).
5 Code and Data Availability
Code and data are available at https://github.com/Arif-PhyChem/txl_fusion
{suppinfo}
Comprehensive methodology and supplementary analyses—including feature engineering, TXL Fusion architecture and training details, a semantic versus numerical descriptor performance comparison, element-resolved performance and confidence analysis with dedicated discussions of low-complexity compounds and chemical sparsity generalization, DFT computational settings, and an assessment of noise and reliability in DFT-based topological labels—are provided in the Supporting Information (Sections S1–S7).